8-pack control system working group Meeting Notes

29 April 2003


In attendance: R. Chestnut, J. Frisch, K. Kim, J. Nelson, J. Rock, S. Smith

Note taker/typer: J. Nelson

Changes in blue.



-Steve reported that later this week we will have 1 PHIT installed which will allow us to read 16 ADC channels.  More PHITs come in a month or so.

-(no change) On going: Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots.



-nothing new



DAC/ADC Issues

-(RC/KK) to do: Some displays need a precision of 5 to get gains to show correctly.


All of these issues are for later:


-TBD: The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-TBD: the trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-Kukhee reported that the following is done and OKOK:

Kukhee reported that Andrew requested some channel and polarity swaps for the DIO.  The swaps are done.  Need to test.


-(no change) 8-pack backup hardware trip circuit 

Marc says yes.  NLCTA ops will take care of it.  Will need SW support for a fault counter, status and reset.  Marc thinks we’ll use Joe’s interlock at least for now.


-This is on hold for now:  SWE will make a PV for a temporary DIO channel as a reset to Joe Frisch’s kludge RE Interlock.  They will send the name to Janice who will put a button on the panel with the pulse width and the klystron attenuation.  Its DIO channel assignment should be on Andrew’s previously sent list.  Also need a channel for status when it trips off. 


-Still to do: From Ron for Judy: check for one in and one out available on DIO for the HW backup trip circuit.

Added to this: Judy will bring a printout of the DIO channel assignments to the next meeting.

Also, Joe reported that we will want more DIO channels for other trip signals, namely those from the modulator and the diodes.

Is the DIO module # the same as the DIO for ATF that we can steal it?  If we need more channels.  If so, need ch assignments and name and the processing loop and panel needs to change.


-(no change) The following item was deferred for a few weeks:

LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-(no change) LLRF folks need to check mini-dbedit files for TRBRs and report problems if any



-Joe reported that we need two more SISes, to be installed in TR09, for modulator signals.  (Assume there will be 10 eventually)

      Ron will talk to Kristi about addresses for these modules (as well as for the SISes to be installed in TR00)

      Joe will provide names and channel assignments – the first SIS module will be 8 modulator signals, the 2nd will be 8 AES signals

      Janice will think about a support panel

      Need to have these 16 signals archived on a fault

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?

-(no change) Right now the SISs are paused and read out with a user request only.  A read when a fault happens will also be added to this.



-Sooner than later Janice will check current list of archived variables to see if any are missing.  (Notably PAVE, PMAX, PMIN, PWR and the processing control).

-(no change) Status? Bob reported the compiled version works.  Instructions on the web. 


Archiving of SISs

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)



-nothing new 



-KK and SS will work to move this to production:  Kukhee made the required database changes, which need to be moved to prod (i.e. reboot TR08)

-Judy made panels for each IQA to enter in these numbers.  Steve and Janice will look.

-Still to do (maybe Judy send Bob updated list?): For Bob: These 6 (or 7) numbers x 8 channels x 4 IQAs need to be archived at least once every 10 minutes. 


Processing algorithm

- Steve confirmed that LLRF folks will indeed contact SWE:  (still waiting on HW folks) Processing algorithm ready to be checked.  Andy will call Kukhee and Ron when there’s a good time to check it.

- Kukhee will make sure the processing loop resets the IQAs too.

-The trigger stopping mechanism on a fault was discussed.  Joe proposed a DIO bit goes to NIM logic etc etc.  It’s believed that Kukhee’s method of changing the trigger register status is too slow for the number of SISes and their location in different crates.


Peak Power Meter (PPM)

-This work continues to slide, but Janice will try to up its priority.

-Tonee and Janice will hook up enet box(es) and gpib cables to the PPMs.

-(no change) Tonee will tweak the labview program to send the waveform to epics once we have names – probably late next week

-(no change) Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)

-Janice will make a panel or two (or at least design them) and teach llrf how to get at archived data and panel.

These PVs will also need to be archived…

Steve will talk to Bobby about perhaps a more permanent home for the PPM besides the stool by klystron 8.



-DCM updates are flaky.  Do other DCMs on site work?  Ron will ask Steph.  He also mentioned there might be an upgrade in our future to bypass the DCM via Ethernet.

Janice will email Alan about the ongoing situation.

-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-many printer questions that led to discussion of rebooting the CmdSvr.  Does someone just need to make an executive decision when gtw04 and 05 can have their cmdsvr’s rebooted?

Is this resolved?  Steve will try to use Ron’s workaround for getting epics to printout on the color printer near his office.  Either with the vax queue name (SLC:Qname) or the NT queue name (something similar).

-Steve will talk to Dave Schultz about getting an NT box for the dance floor.

-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.

-Has this been done?  Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman

-(no change) we deferred the following:

-LLRF needs to decide what the reset button does.


We also discussed the need for spy cameras on the dance floor.  Sounded like Joe was going to look into it.